OpenStudy (anonymous):

the sum of two consecutive even integers a and b is 5/12. find integers

OpenStudy (mathstudent55):

How can the sum of two integers equal a fraction?

OpenStudy (anonymous):

since they are consecutive we have b=a+2 and since the sum is 5/12 v have a+b=5/12 solve these equations to get ur result

OpenStudy (anonymous):

ops u r right @mathstudent55 i didnt notice this point

OpenStudy (anonymous):

??

OpenStudy (anonymous):

r u sure it is 5/12? @momma821

OpenStudy (anonymous):

yes im sure

OpenStudy (anonymous):

there are no two even integers whose sum is 5/12, if u still wana try, u solve the two equations i made u will get some values of a and b but they wont b even integers.

OpenStudy (anonymous):

ok thanks!!

OpenStudy (anonymous):

np